OpenStudy (anonymous):

What is the domain of f(x)? A. -2 < x ≤ 3 B. -2 ≤ x < 3 C. -1 < y ≤ 3 D. -1 ≤ y < 3

OpenStudy (solomonzelman):

what is the f(x), lol?

OpenStudy (secret-ninja):

Is that the whole question?

OpenStudy (solomonzelman):

Well, you can exclude C and D, because y is the range, not the domain:)

OpenStudy (anonymous):

Hold on theres a graph

OpenStudy (anonymous):

The chat box is on the way of the "post" button

OpenStudy (secret-ninja):

Just close the chat box.

OpenStudy (solomonzelman):

click Tab+enter, if your post button is hiding.

OpenStudy (phi):

or refresh your browser

OpenStudy (anonymous):

Thank you guys

OpenStudy (solomonzelman):

Open dot = the point is not included. Closed dot=the point is included.

OpenStudy (anonymous):

I put B

OpenStudy (anonymous):

But idk if i'm correct

OpenStudy (solomonzelman):

Yes, I wouldn't say you are correct. Do you understand what I mean by my last (before this) post?

OpenStudy (anonymous):

huh ?

OpenStudy (solomonzelman):

``` Open dot = the point is not included. Closed dot=the point is included. ```

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

So that would mean A ?

OpenStudy (solomonzelman):

yes, duh, after I said B is incorrect and C and D are excluded. Critical thinker, lol

OpenStudy (anonymous):

lol okokok Thank you

OpenStudy (solomonzelman):

Yw

OpenStudy (anonymous):

The equation of a line is shown below: y = - x - 7 What is the equation of a line which is parallel to this line and passes through (1, 79)? A. y = -x + 80 B. y = x - 80 C. y = 7x - 80 D. y = -7x + 80

OpenStudy (anonymous):

Got D

OpenStudy (solomonzelman):

what are you chainging here, the slope or the y-intercept?

OpenStudy (solomonzelman):

*changing

OpenStudy (anonymous):

But i couldn't really understand this one

OpenStudy (anonymous):

The slope

OpenStudy (solomonzelman):

What difference do you see between the 2 lines?

OpenStudy (anonymous):

intersection ?

OpenStudy (anonymous):

irdk